# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

### v1

@@ 4. Results
4.1. RQ1-Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in individual and group settings?
Participants received a collection of 36 explanations to solve the study tasks and decide about the system’s deployment individually and (in focus groups) collectively (Section 3). The following subsections describe the explanations’ effect on participants’ understanding in both settings (4.1.1), design aspects that impeded understanding (4.1.2), the time limit’s effect on prioritization of information (4.1.3), the interaction differences between single and group settings (4.1.4), and participants’ ratings of explanations’ helpfulness and influence (4.1.5).
4.1.1. Explanations had mixed effects on participants’ understanding.
| Focus gropus | ||||||||||||||
| Increased understanding | F2 | C3 | F1 | B4 | A2 | H1 | H2 | E1 | H4 | |||||
| Self-reported understanding | +++ | ++ | ++ | + | + | + | + | + | + | |||||
| Decision confidence | - | ++ | = | +++ | + | + | + | = | = | |||||
| Unchanged understanding | H3 | B3 | A3 | G1 | H5 | A1 | A4 | C2 | D3 | E2 | C1 | E3 | F4 | F5 |
| Self-reported understanding | = | = | = | = | = | = | = | = | = | = | = | = | = | = |
| Decision confidence | +++ | ++ | + | + | + | = | = | = | = | = | - | - | - | - |
| Decreased understanding | D1 | G4 | D2 | G2 | B1 | B2 | F3 | G3 | ||||||
| Self-reported understanding | - | - | - | - | - - | - - | - - - | - - - | ||||||
| Decision confidence | + | + | = | = | +++ | = | + | = | ||||||
| Single studies | S6 | S12 | S8 | S7 | S4 | S9 | S1 | S3 | S11 | S5 | S10 | S2 | ||
| Self-reported understanding | +++ | + | + | = | = | = | = | = | = | - | - | - - | ||
| Decision confidence | ++ | + | = | +++ | + | + | = | - - | - - | ++ | = | - |
| Focus groups | Single interviews | |||||||||
| ID | 1st decision | Collective decision | 2nd decision | ID | 1st decision | Collective decision | 2nd decision | ID | 1st decision | 2nd decision |
| A1 | No | Yes | F1 | Yes | Yes | S1 | Yes | Yes | ||
| A2 | No | Yes | F2 | Yes | Yes | S2 | No | No | ||
| A3 | Yes | Yes | F3 | Yes | No | S3 | No | No | ||
| A4 | No | Yes | Yes | F4 | No | No | S4 | No | No | |
| B1 | No | No | F5 | Yes | No | Yes | S5 | Yes | No | |
| B2 | Yes | No | G1 | Yes | Yes | S6 | No | No | ||
| B3 | Yes | No | G2 | No | No | S7 | No | No | ||
| B4 | Yes | No | No | G3 | Yes | No | S8 | No | No | |
| C1 | No | No | G4 | No | No | No | S9 | No | No | |
| C2 | No | No | H1 | No | No | S10 | Yes | Yes | ||
| C3 | No | No | Yes | H2 | No | No | S11 | Yes | No | |
| D1 | No | No | H3 | No | No | S12 | No | No | ||
| D2 | No | No | H4 | No | No | |||||
| D3 | No | No | Yes | H5 | No | No | No | |||
| E1 | No | No | ||||||||
| E2 | No | No | ||||||||
| E3 | No | No | Yes | |||||||
Participants reported their understanding before and after receiving and discussing the explanations. Table 3 shows that self-reported understanding decreased and increased for comparable numbers of participants, while also often remaining unchanged. Participants gave divergent feedback on how they perceived the explanations, describing their structure and design as ”nicely presented” (A2, C2), ”good to get an overview” (C3, H4), ”well-selected” regarding the questions (G2), ”active and controllable” (S8), and as a good primer (F4). Critical comments described them as ”too much” (D1, S4), ”at first overwhelming” (H5), ”confusing” (B1, D1), ”theoretical” (D2), and ”quite complex” (C2). However, this was not always seen as an issue: ”I believe that it is complex. But I also believe that it needs to be.” (D2).
Participants gave nuanced reasons why their perceived understanding decreased. Several focus group participants stated they did not fully understand the technical explanations, including numbers and a formula (B3, C3, G4). In contrast, participants in single interviews with comparable backgrounds met fewer challenges with technical explanations, often stating that they appreciated applying this knowledge when solving the tasks (S2, S8, S11). In these cases, the setting likely mediated the ease of interaction, as participants in single interviews could concentrate without distraction (further described in Section 4.1.4+ 5.2). However, irrespective of the setting, confrontation with mathematical details seemed to re-adjust participants’ benchmarks for ”high understanding” and calibrate their perceptions. Likewise, several participants stated that the different subtopics and levels of detail in the explanations helped them identify the information they were still missing to judge the whole system (B1, F3, G3). Both observations indicate that the explanations revealed participants’ blind spots and helped them to overcome the ”illusion of explanatory depth” (Rozenblit and Keil, 2002) – an overestimation of one’s understanding due to limited knowledge and misled intuition – which is exemplified by S2’s comment: ”The more I look at the explanations now, the more I realize that I simply assumed a lot of things but didn’t actually know how it worked.” In short, participants gauged their understanding relative to the amount of information available. That is, since participants only had time to read through parts of all available explanations and were aware that they had not seen all of them, they reported a decrease in perceived understanding when it, in fact, improved. The implications for measuring explanations’ effect on understanding are discussed in Section 5.
4.1.2. The explanation design was not equally suited to every participant group.
In contrast to participants whose perceived understanding was improved and calibrated by the explanations, we observed that the explanations did not fully support understanding or interaction for some groups. In these cases, the explanations’ language and textual format were listed as impeding factors. The language style was described discrepantly as ”well understandable” (P3, E3, H1) but also ”very difficult” (E2, F2) and ”too technical” (B3, B4) depending on participants’ background. Crucially, in Groups E and F (decision subjects), participants stated that the language posed challenges. E2 compared the explanations to official letters they receive: ”I get letters that I don’t understand from the court. I understand every single word, but I don’t understand the context.” In line with this, job counselors in Groups C, G, and H stated that explanations in simple language would be required if they were supposed to be useful for their job-seeking clients. Similar comments were made about the explanations’ heavy reliance on textual information, described as demanding and requiring plenty of concentration. While previous work found that textual explanations can be effective despite users’ dispreference (Szymanski et al., 2021) or are even preferred (Weitz et al., 2024), a more interactive and visual design would likely facilitate interaction: ”If there’s a program where, when you click on it, it opens as a window and you can read about it, that would be great, and you can deal with it at your own pace.” (H5) Both, simpler language options and an interactive version, are thus useful suggestions for improving the explanation design to ensure that explanations are helpful for a broader range of stakeholder groups. Further, Section 4.2.1 describes how these understanding challenges impacted group interaction.
4.1.3. Time limits led to forced prioritization of information, with different effects on participants.
All participants had 40 minutes to become familiar with the explanation structure, request new explanations, solve the given tasks, and, in focus groups, make a collective decision (Table 4). The time limit was meant to incentivize participants to prioritize information and, where possible, work together to process larger amounts. This succeeded for most focus group’s participants, who stated that ”when you have little time, the teamwork is extremely good” (B2) and that ”even a brief exchange is enough, otherwise it becomes too much” (C3). However, the time limit also led to induced stress, especially in single interviews (S2, S6, S7, S8, S9, S11). S9 commented that this reminded them of their work as a social center manager: ”Basically, the time to read about something new does not exist.” (S9). This points to the need for brief and compact synopses complementing the detailed explanations. Regarding prioritization, many participants liked choosing explanations according to their own interest, which gave the feeling of being in control (S8), provided a direction for exploration (H1), and justified skipping explanation categories in favor of others (S6, S9). Allowing for the self-directed selection of relevant information by users could thus be beneficial to encourage interaction and engagement.
4.1.4. Some single participants worked through surprising amounts of explanations.
Compared to focus groups, participants in single interviews often requested and engaged with equal or even higher amounts of explanations (as depicted in Figure 4). Often, participants in single interviews requested ”the whole context” (S3) or ”all of level 3” (S4) of one category, appearing to deal with the amount of information easily. A likely reason for this is that single interviews supported more focused and in-depth interaction with explanations compared to group settings. Participants in single interviews also performed better in the study tasks overall and, as mentioned above, often calculated the job-seeker’s precise employment chance in task 2, which required following a multi-step process. This is noteworthy, as none of the groups, aside from the pilot groups, attempted to calculate the precise chance. This cannot be explained by a difference in educational or professional background, as even focus groups with university graduates in domain expert professions did not perform these calculations, again indicating that the reason might be the difference in supported interaction. For example, the parallel processing of information might have impeded understanding of the system’s entirety in groups, as described by G3: ”You only pick out the most important things. And then you haven’t gathered information about the thing as a whole.” This highlights the benefits and trade-offs of both settings, suggesting that individual engagement with explanations might be helpful before discussing them in a group setting. The implications of this are discussed further in Section 5
4.1.5. All explanation categories were deemed helpful and influential, but system details and context were most often mentioned.
In the second self-reports, participants were asked which explanation category was most helpful for their understanding, most influential for their decisions, and whether specific explanations stood out. Regarding understanding, participants in focus groups responded with a rather even distribution over all categories, as depicted in Figure 5, and frequently mentioned that ”all of them [are relevant]… I don’t think you can leave anything out, really” (D3). Both participants in focus groups and single interviews further stated that system details were crucial, as they were perceived as ”tangible” (S6), ”concrete” (S8), and to be ”the devil in the detail” (B4). Specifically, participants found explanations about the features and weighting most important (D1, G1, H3, S5), which G4 commented with: ”That is the central point, the basis of the whole system.” Even though context was less often reported as an influential explanation category, participants requested these explanations most often (Figure 6) and frequently mentioned affected persons’ inability to contest decisions (C1, D1, E1) and the system’s political background (A1, S3) as relevant – both belonging to context. Notably, participants in single interviews found data much less helpful and less influential, stating, e.g., that GDPR would likely cover the important aspects of it (S8) or that they prioritized another category in the time available: ”I didn’t deal with it in such detail […]. You read it [the base level] and already feel informed. As is often the case with data protection, ’yeah, close enough’.” (S6). Other participants stated that the technical details were not too important for them but that they instead focused on how the system was used practically and which values it represented (S3, S10), corresponding to the categories usage and context.
4.2. RQ2-Deliberation: How do the explanations support group interaction with information and deliberation processes?
In this section, we describe how our explanations supported group collaboration in exchanging information (Section 4.2.1) and in discussing and forming opinions about the system’s deployment (Section 4.2.2).
4.2.1. Explanations support collaborative success if group dynamic is constructive.
We identify several processes that groups used to interact with information collectively. Our descriptions of these interaction processes include splitting (by category, order, priority), sharing (swapping, reading out loud, proposing solutions), analyzing (discussing, explaining), and combining (merging, locating, correcting). The degree to which groups used these processes appeared to depend on the familiarity between group members, their domain expertise, and their understanding. In the following, we analyze three exemplary conversation excerpts that showcase different forms of interaction in focus groups while also demonstrating the diversity in collective interaction. For each excerpt, we describe the interaction processes that emerged and then connect them to mechanisms that drive collaborative failure or success (marked by quotation marks) as outlined by Nokes-Malach et al. (2015).
Case 1: Information sharing leads to increased engagement and construction of common ground. Group G was composed of participants who hold executive positions in a civil society organization and are responsible for the management and training of employees. Participants in this group showed high engagement with information, frequently sharing explanations and exchanging their perspectives by tapping on their professional experience. The excerpt begins several minutes after participants received the explanations just as G4 examines an explanation comparing the algorithmic score for two job-seekers:
@@
G4: That’s bad, the two of them. Look, I’ve got this [system details-C-base], ”What chances would the system calculate for Martin and Schifteh?” (sharing: reading out loud)
G2: Schifteh is probably worse off, isn’t she? (analyzing: discussing)
@@
G4: Schifteh has 30% chance of employment and Martin 52%, even though Schifteh has a degree and would be working in the IT sector. And Martin has compulsory schooling and works in the cleaning sector. Martin’s chances of employment are almost twice as high as Schifteh’s. (sharing: reading out loud; analyzing: explaining) […]

### v2

@@ 4. Results
4.1. RQ1-Explanations: How does a question-driven, modular explanation design support AI novices’ understanding in groups and individual settings?
To examine how AI novices used the explanations to understand the study’s use case, we analyzed their self-reports, articulations, and interactions in both settings. We found that each setting supported different aspects of understanding, suggesting a trade-off. We first describe how the explanations contributed to shared understanding and ‘collaborative success’ in groups (4.1.1) and continue with the explanations’ role in instances of ‘collaborative failure’ (Nokes-Malach et al., 2015) (4.1.2), summarized in Figure 5. We then describe individuals’ interactions with the explanations (4.1.3), participants’ feedback on the explanation design (4.1.4), and summarize the benefits and drawbacks of both settings for XAI (4.1.5).
4.1.1. Groups’ benefits: Shared understanding and increased engagement.
In the best cases, groups leveraged the modular explanation structure to use distributed cognition (Keil, 2003), meaning that participants processed information in parallel and then combined it. We use the term shared understanding to capture interactions that realized distributed cognition. Examples of such interactions included locating information together, sharing information with others, discussing interpretations, debating task solutions, and querying and explaining (a question by a group participant invites other participants to contribute). The explanations only afforded this set of interactions to groups, as they required social interaction with other participants. For example, in Group C, participant C1 read the first study task aloud and asked for input (querying), after which the group discussed solutions (explaining):
C1 Can Harald change the data stored about him? Yes, he can certainly change it, can’t he? […]
C3 Which stored data, the one down there? [points at Harald’s demographic features]
C1 Yes, just that.
C3 49 – no, male – no. The apprenticeship – no, Austria – he can still change that. Duty of care – he could get married or have children. He could change his service sector. He could change his career. Impairment…
C1 Well, what is meant by ‘change’? When he enters the data, he can change the data. He doesn’t have to specify the knee problem. […]
C3 So he can change it.
C2 Yes.
Interactions such as querying and explaining and discussing interpretations rely on collaboration between participants to ‘share working memory resources’, ‘complement others’ knowledge’, ‘re-expose information’, and ‘correct errors’. These aspects of collaboration are described as cognitive mechanisms of ‘collaborative success’ (Nokes-Malach et al., 2015) and provide groups with multiple ways to tackle explanations. For example, participants tended to work through information about usage and context alone or in pairs but raised explanations with the group when they were difficult or piqued their interest. This was often the case with explanations about system details, which included the most numerical information but also were an important key to understanding the system’s calculations. We describe the process of using other’s understanding to close gaps in one’s own as outsourcing (Keil, 2006). As understanding AI systems involves interacting with a variety of different information categories (e.g., technical, political, social), outsourcing provides a way to hand information to the team member most competent in this category. For example, in Group B, B2 expressed their appreciation for B3’s help in solving study task 2: “It was a math problem. You [B3] filtered it out well. It was very analytical. With your help, we were able to recognize these weak points.” In contrast to querying and explaining, which participants used to invite input or spark conversation, outsourcing was thus used for the active delegation of an impeded understanding process.
Whether groups used this collaboration depended on participants’ relationships and the group’s social dynamics. The explanations also served to support these ‘social mechanisms’ of collaboration, by drawing the group’s attention to certain aspects of the AI system and encouraging them to share their experiences and opinions. We present an excerpt from Group G as an illustration. Here, participant G4 shared an explanation that documented the algorithm’s impact on two job-seekers (‘joint management of attention’), which prompted G2 and G3 to discuss interpretations (‘increased engagement’). This interaction established ‘common ground’ that the group later used for deliberation:
@@
G4 That’s bad, the two of them. Look, “What chances would the system calculate for Martin and Schifteh?”
G2 Schifteh is probably worse off, isn’t she?
@@
G4 Schifteh has a 30% chance of employment and Martin 52%, even though Schifteh has a degree and would be working in the IT sector. And Martin has compulsory schooling and works in the cleaning sector. Martin’s chances of employment are almost twice as high as Schifteh’s. […]